Parth (parthkohli):

I am a basic trigonometry learner. I have heard till now that trigonometry is all about right triangles. Please help.

Parth (parthkohli):

http://assets.openstudy.com/updates/attachments/4fcb45c2e4b0c6963ad4f399-mandy17-1338721738166-71930792159341b491eb279077d0b514.gif I don't understand how we can get the required side lengths when this isn't a right angled triangle.

Parth (parthkohli):

I mean, how can even trig have such things when the fundamental is to use right triangles.

Parth (parthkohli):

I just need to know how it is used in non-right triangles. I need not find the other lengths.

OpenStudy (anonymous):

Many different tools. You have the fundamental Law of Cosines, Law of Sines, and the fact that all the angles of a triangle add up to \(180\) degrees.

Parth (parthkohli):

I mean that this isn't used in right angled triangles. How can trig be without right triangles? o.O

OpenStudy (anonymous):

Trigonometry is a general study of triangles. Only a subset of trigonometry is right triangle trigonometry.

OpenStudy (asnaseer):

The sine/cosine rules are derived from right angled triangles. You can see a proof of the sine rule here if you are interested: http://www.youtube.com/watch?v=DsSVxuFh1WA

OpenStudy (anonymous):

Get the third angle , Use the Law \[/\huge \frac{a}{sinA} = \frac{b}{sinB} = \frac{c}{sinC}\]

Parth (parthkohli):

Hmmm....I see.

OpenStudy (anonymous):

Wiki has a nicer definition of trigonometry: "Trigonometry (from Greek trigōnon "triangle" + metron "measure"[1]) is a branch of mathematics that studies triangles and the relationships between their sides and the angles between these sides."

OpenStudy (lgbasallote):

sine law and cosine law are very broad o.O a lot of conditions

Parth (parthkohli):

I just needed to know how and why it's used in non-right triangles. Thank you :)

OpenStudy (anonymous):

You are welcome. :)